### v3

@@ Appendix A Model Properties
The number of parameters in BERT-like models varies significantly depending on their architecture (see Table A). The base version of BERT has approximately 110 million parameters, while the large version expands to 340 million. RoBERTa, an optimized variant with improved pre-training, features 125 million parameters in its base model and 355 million in its large model.
Multilingual models like XLM-RoBERTa are designed for cross-lingual tasks, with the base version containing 270 million parameters and the large version 550 million. Electra, using a generator-discriminator framework, achieves competitive performance with fewer parameters: around 110 million in the base model and 335 million in the large model.
For German NLP, GeistBERT introduces several architectural variants optimized for efficiency and extended context. The standard GeistBERT model contains approximately 125 million parameters, matching the configuration of GottBERTbase. The GeistBERTnyströmformer variant integrates sparse Nyström attention to support longer input sequences and improve scalability, resulting in a slightly larger model with 131 million parameters. Meanwhile, GeistBERTlongformer uses sliding-window attention to enable extended context lengths and has around 153 million parameters. Other widely used German models, such as GBERT, GELECTRA, and dbmdz BERT, adopt architectures comparable to BERTbase ($\sim$110M) and BERTlarge ($\sim$345M), but differ primarily in their pretraining corpora and learning objectives, rather than being specialized for a specific domain.
@@ Appendix C Parameters
Table 7: Training (TT) and evaluation runtime (ET) in seconds of the best downstream task models for each task. #GPU denotes the number of devices required for training and GA the gradient accumulation steps. The evaluation could be run on a single GPU.
Task GeistBERT GeistBERTnyströmformer GeistBERTlongformer
#GPU GA TT ET #GPU GA TT ET #GPU GA TT ET
XNLI 1 1 47,684 20.0 1 1 83,100 23.9 1 1 33,404 41.2
CoNLL 03 1 1 807 2.2 1 1 891 2.7 2 4 14,611 30.7
GermEval 2018 fine 1 1 200 9.0 1 1 754 4.0 1 1 4437 7.0
coarse 1 1 196 3.6 1 1 979 3.8 1 1 4788 8.8
GermEval 2014 1 1 2,553 2.6 2 1 19908 66.0 2 1 22644 54.8
10kGNAD 1 1 1,219 3.9 1 1 1,552 4.5 2 4 11,159 16.6

### v4

@@ Appendix A Model Properties
Table A lists the vocabulary sizes and total parameter counts of all models included in our evaluation. While most German BERT-style base models, such as GBERTbase, dbmdzBERT, and GELECTRAbase, contain approximately 110 million parameters, GeistBERT and ${}^{\mathrm{f}}\mathrm{GottBERT}_{base}$ are slightly larger at around 126 million parameters due to their RoBERTa-based architecture and a larger vocabulary of 52,009 tokens.
Large-scale German models such as GBERTlarge, GELECTRAlarge, and ${}^{\mathrm{f}}\mathrm{GottBERT}_{large}$ contain between 335 and 357 million parameters. Among the multilingual models, XLM-RoBERTabase and XLM-RoBERTalarge are substantially larger, with 278 million and 560 million parameters respectively. The vocabulary sizes vary across models and are influenced by tokenizer design and pre-training data. GeistBERT uses the same tokenizer as GottBERT, which is based on byte-level BPE trained on German text.